OpenStudy (anonymous):

What is the average rate of change from x = 1 to x = 3? -4 -2 2 4

OpenStudy (anonymous):

@ganeshie8 @satellite73 2??

ganeshie8 (ganeshie8):

average rate of change of whom ?

OpenStudy (anonymous):

why can't i post the graph

OpenStudy (anonymous):

when i attach, it won't let me press the post button

ganeshie8 (ganeshie8):

Oh, refresh the page and try again

OpenStudy (anonymous):

yep that worked

ganeshie8 (ganeshie8):

good :) average rate of change between two points is simply the slope of line that joins them

OpenStudy (anonymous):

ok thanks can you help with 1 more

OpenStudy (anonymous):

y = -1(x - 4)^2 + 1 y = -1(x + 1)^2 + 4 y = 1(x + 1)^2 - 4 y = 1(x + 4)^2 - 1

OpenStudy (anonymous):

what is the equation of the graph I don't know how to do this one

ganeshie8 (ganeshie8):

the parabola is opening UP, so the value of "\(\large a\)" has to be positive in : \[\large y = a(x-h)^2+k\]

ganeshie8 (ganeshie8):

in light of above information, which two options can you eliminate ?

OpenStudy (anonymous):

makes sense, so a and b are out

ganeshie8 (ganeshie8):

Yes! look at the graph, whats the vertex point ?

OpenStudy (anonymous):

(-1,-4)

ganeshie8 (ganeshie8):

perfect! so \(\large h = -1\) and \(\large k = -4\)

OpenStudy (anonymous):

so c?

ganeshie8 (ganeshie8):

Correct!

OpenStudy (anonymous):

ok thanks so much, those were the two i wasn't sure about. you make everything seem easy!

OpenStudy (anonymous):

they were right, thanks again ^_^

ganeshie8 (ganeshie8):

np :) yw !